What is the next electrostatic force of particle
1. Due to particles 2 and 3, If Q1,Q2, and Q3 are 1.6×10^-19 colombs, 3.2×10^-19 Colombs and -3.2×10^-19 colombs. Q3 is at a distance ¾R from Q1.
1
Expert's answer
2019-08-29T09:42:54-0400
"F_{12}=-(9\cdot10^9)\frac{(1.6\cdot10^{-19})(3.2\cdot10^{-19})}{0.02^2}=-1.15\cdot10^{-24}N"

"F_{13}=(9\cdot10^9)\frac{(1.6\cdot10^{-19})(3.2\cdot10^{-19})}{(0.02\cdot 0.75)^2}=2.05\cdot10^{-24}N"

"F=-1.15\cdot10^{-24}+2.05\cdot10^{-24}=9.00\cdot10^{-25}N"